# Instruction Fine-Tuning: Does Prompt Loss Matter?

### v3

@@ Abstract
Mathew Huerta-Enochian  and Seung Yong Ko EQ4ALL 11 Nonhyeon-ro 76-gil, Gangnam-gu, Seoul {mathew, stephenko}@eq4all.co.kr
@@ 1 Introduction
While training, model parameters are updated by optimizing for next-token maximal likelihood classification. Most open sourced solutions for SFT either mask the prompt loss (for prefix language modeling) or use the entire sequence loss (for full language modeling) while some API providers support an explicit PLW parameter that allows users to apply fractional PLW during SIFT. The commonly-held notion is that fractional PLW helps stabilize when fine-tuning on data with short outputs. Recently, however, OpenAI quietly removed support for their prompt_loss_weight parameter.
The reason for the removal is unknown. Furthermore, to our knowledge, there has never been a proper study on the effects of PLW.
@@ 2.1 Definitions
We define the generation ratio $R_{g}$ as the ratio of completion length to prompt length. We then divide instruction data into two broad categories. Data with $R_{g}<1$ are short-completion data, and data with $R_{g}\geq 1$ are long-completion data. When applied to an entire dataset, we take $R_{g}$ to be the mean completion-prompt ratio.
@@ 3 Hypotheses
Based on these assumptions, we would expect the two competing factors to result in a downward curved relationship between PLW and downstream performance. Limiting PLW to the range of [0, 1], we postulate that there is a critical value $\lambda$ for PLW with $10<=\lambda<=$ 1. For PLW less than $\lambda$, the positive effect dominates the negative effect and for values greater than $\lambda$, the negative effect dominates the positive effect. If $\lambda=0$, then PLW’s contribution to model performance is strictly negative, and if $\lambda=1$, then PLW contributes strictly positively to model performance. Note that $\lambda$ is not an intrinsic characteristic of the dataset, model architecture, or training algorithm. Rather, it would depend on numerous factors and change for each task.
@@ 5.1 Performance Trends
A group of four benchmarks (Arc Challenge, PIQA, TruthfulQA-Gen, and WinoGrande) clearly show the expected negative quadratic relationship between wp and performance of AlpacaDataShort models. Notably, AlpacaDataShort models outperform AlpacaData and AlpacaDataCleaned on this group given optimal PLW tuning. On the long-generation benchmarks, however, AlpacaDataShort models show a steadily-increasing trend. For these seven benchmarks and AlpacaDataShort models, using prompt loss masking (wp$=0$) almost always produced the worst scores.
The wp-based performance increase was around twenty percentage points for both long-generation benchmarks. Contrasted with the less than two percentage point change for short-generation and multiple choice benchmarks, and it is evident that the optimal PLW for short-generation and long-generation benchmarks is different.
Clear qualitative performance trends for the AlpacaData and AlpacaDataCleaned and for any model on the six translation benchmarks could not be identified.
@@ 5.3 Causal Mechanism & Interpretation
We next explored how PLW affected training data memorization. We sampled 10,000 unique prompts from the AlpacaDataShort training set, generated completions from each prompt, and calculated corpus BLEU-4 scores.
We found that for PLW from 0.0 to around 0.1, models memorized most of the training data, consistently scoring near 80 corpus BLEU. Corpus BLEU then decreased as PLW increased from 0.1. We also analyzed generation length on the AlpacaEval 1 test set which showed high deviations in generation length for the high loss RSD PLW values for LLaMA 2 only and a generally increasing length trend with PLW. Considering that AlpacaDataShort contains short completions, this strongly suggests that PLW contributes a regularizing effect and helps decrease overfitting.
The positive relationship between performance and PLW on AlpacaEval 1 and PandaLM as opposed to the negative quadratic relationship seen on several of the other benchmarks can then be explained by a tradeoff between longer generations and decreased overfitting for high PLW and better preservation of PTLM weights for low PLW.
@@ 6 Supplemental Experiments
In this section, we present supplemental PLW experiments that suggest that PLW cannot be replaced by alternative regularization techniques for SIFT and that the effects of PLW extend to other short-completion datasets. Since the translation benchmarks showed high levels of noise and unclear correlation with PLW in the main experiment, they were not included in supplemental experiments.
@@ 6.1 Alternative Regularizers
To investigate if the effects of PLW on short-completion SIFT can be emulated with other common regularization techniques, we repeated the AlpacaDataShort training runs for PLW $=0$ and PLW $=1$ while applying various regularizers. We tested weight decay, minimizing the Minkowski distance between PTLM weights and learned weights, dropout, and label smoothing. See Appendix C.1 for additional visualizations.
@@ 6.2 Alternative Datasets
We repeated our main training experiment using two additional datasets, a cleaned and binarized version of UltraFeedback (Cui et al. (2023); denoted UltraFeedbackCleaned) and databricks-dolly-15k (Conover et al., 2023; denoted DatabricksDolly). We also trained on prompt-inverted versions of both datasets, denoted UltraFeedbackShort and DatabricksDollyShort, respectively, and expanded analysis to include LLaMA 3 8B AI@Meta (2024) as a PTLM. See length statistics for these four datasets in table 1 and appendix C.2 for additional visualizations.
The combined relative aggregate scores for UltraFeedbackCleaned and DatabricksDolly and for UltraFeedbackShort and DatabricksDollyShort are visualized in figure 3. Visual inspection suggests that PLW affected learning for both groups of datasets. For the shortened variants, performance appears to have the same negative quadratic relationship with PLW as AlpacaDataShort models do in the main experiment. The relationship between PLW and performance for models fine-tuned on the original datasets is less clear but appears to have a generally increasing trend, unlike the long-completion datasets in the main experiment. We believe this is due to both UltraFeedbackCleaned and DatabricksDolly having much smaller completion-prompt length ratios $R_{g}$ than the AlpacaData and AlpacaDataCleaned datasets.
@@ 7 Conclusion
We found that PLW has a statistically significant effect on learning for our short-completion dataset, and proper tuning of PLW allowed short-completion-trained models to outperform long-completion-trained models on short-generation benchmarks. We showed that the causal mechanism was due to a balance between two different regularizing effects and not due to increased training stability as is commonly attributed. We showed that the measured relationship extends to additional SIFT datasets and that the effects could not be sufficiently emulated with other regularizers.
@@
    Given the importance of PLW and given that many SIFT datasets and almost all natural language understanding (NLU) datasets are short-completion datasets, we warn SFT API providers about the need for a PLW parameter to adequately cover a full range of modeling applications.
@@
    We analyzed prompt loss weighting (PLW) for instruction fine-tuning LLMs. We characterized three fine-tuning datasets by their relative completion-prompt length ratios and reported on the effect of PLW when training on each dataset. It would be helpful to extend this research to a wider range of datasets to increase the strength of our conclusions and create more complete guidelines for prompt loss weighting.
@@
    LLM-as-evaluator approaches like PandaLM and AlpacaEval 1 are still relatively new, and these approaches are being actively developed. We chose to use Mixtral 8x7B as an auto-evaluator for AlpacaEval 1 due to budget limitations. While we cite high human evaluation correlation with Mixtral and justify this decision in section 4.4, using the default auto-evaluator would be beneficial for better comparison with other research.
@@ Acknowledgements
This work was supported by the Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City (No. BA00000797, LLM-based sign language translation for weather forecasts).
This research was possible thanks to the support of our coworkers at EQ4ALL.
@@ Appendix A Prompt Inversion
In order to experiment with short-completion fine-tuning datasets while also preserving a similar dataset to use as a reference, we propose using prompt inversion to rotate the instruction, input, and output fields of instances. Prompt inversion modifies an instance to use the original instruction as the new completion text and the original output as the new input text. The model is then given the following instruction:
@@
To generate short versions of our test instruction datasets, we use prompt inversion to modify every instance with a completion-prompt length ratio $R_{g}>1$, based on the tokenized lengths of each field and using prompt to mean the concatenation of the instruction and input fields as explained in section 2.1.
Synthesis of the original input and output texts to predict the original instruction requires both language understanding and reasoning. It also does not train the model for long-generation (since new completions will generally be short) and thus does not prepare the model for train-of-thought generation that long-generation tasks intrinsically learn. Therefore, we propose that prompt-inverted instances are best suited for learning natural language understanding (NLU)-related tasks rather than general generation tasks.
@@ Appendix E Reproducibility
This section provides technical details on all experiments and benchmarks for transparency and to encourage reproduction of results. To help with reproducibility, we will also upload the fine-tuned models, test generation outputs, and our modified datasets to xx.yy.zz.
@@ E.1 Model Fine-Tuning
Model fine-tuning was performed with the Stanford Alpaca GitHub repository at www.github.com/tatsu-lab/stanford_alpaca/tree/761dc5b.
@@
AlpacaData is available from the Stanford Alpaca repository. AlpacaDataCleaned can be found at www.github.com/gururise/AlpacaDataCleaned/tree/791174f and is labeled “alpaca_data_cleaned.json”. As noted above, we released AlpacaDataShort at xx.yy.zz.
@@ E.2 Model Evaluation
In an effort to match the current HuggingFace Open LLM leaderboard, we evaluated ARC Challenge, TruthfulQA-MC2, WinoGrande, and PIQA on the same EEH commit that the HuggingFace leaderboard uses: www.github.com/EleutherAI/lm-evaluation-harness/tree/b281b09 We also matched the number of shots with the number used for the HuggingFace leaderboard for ARC Challenge, TruthfulQA-MC2, and WinoGrande.
TruthfulQA-Gen and all translation tasks were evaluated using a more recent commit at www.github.com/EleutherAI/lm-evaluation-harness/tree/b93c3bc. We modified the translation tasks at this commit to include an appropriate prompt to support zero-shot translation. These changes can be seen at xx.yy.zz.
@@ E.5 Supplemental Experiments
The second experiment introduced two new datasets: UltraFeedbackBinarizedCleaned and DatabricksDolly. UltraFeedbackBinarizedCleaned can be found at https://huggingface.co/datasets/allenai/ultrafeedback_binarized_cleaned/tree/f304ce5. And DatabricksDolly can be found at https://huggingface.co/datasets/databricks/databricks-dolly-15k/tree/bdd27f4.

### v4

@@ Abstract
Mathew Huerta-Enochian1thanks: Currently affiliated with the University of Oregon, research conducted while working at EQ4ALL.  and Seung Yong Ko2 EQ4ALL 11 Nonhyeon-ro 76-gil, Gangnam-gu, Seoul 1mhuerate@uoregon.edu 2stephenko@eq4all.co.kr
@@ 1 Introduction
While training, model parameters are updated by optimizing for next-token maximal likelihood classification. Most open sourced solutions for SIFT either mask the prompt loss (for prefix language modeling) or use the entire sequence loss (for full language modeling) while some API providers support an explicit PLW parameter that allows users to apply fractional PLW during SIFT. The commonly-held notion is that fractional PLW helps stabilize learning when fine-tuning on data with short outputs. Recently, however, OpenAI quietly removed support for their prompt_loss_weight parameter.
The reason for the removal is unknown. Furthermore, to our knowledge, there has never been a proper study on the effects of PLW. [^1] Note that this research used a general implementation of PLW. OpenAI’s former prompt_loss_weight implementation could not be tested directly.
@@ 2.1 Definitions
We define the generation ratio $R_{g}$ as the ratio of completion length to prompt length (also referred to as the completion-prompt ratio). We then divide instruction data into two broad categories. Data with $R_{g}<1$ are short-completion data, and data with $R_{g}\geq 1$ are long-completion data. When applied to an entire dataset, we take $R_{g}$ to be the mean completion-prompt ratio.
@@ 3 Hypotheses
Based on these assumptions, we would expect the two competing factors to result in a downward curved relationship between PLW and downstream performance. Limiting PLW to the range of [0, 1], we postulate that there is a critical value $\lambda$ for PLW with $10<=\lambda<=$ 1. For PLW less than $\lambda$, the positive effect dominates the negative effect and for values greater than $\lambda$, the negative effect dominates the positive effect. If $\lambda=0$, then PLW’s contribution to model performance is strictly negative, and if $\lambda=1$, then PLW contributes strictly positively to model performance. Note that $\lambda$ would not be an intrinsic characteristic of the dataset, model architecture, or training algorithm. Rather, it would depend on numerous factors and change for each task.
@@ 5.1 Performance Trends
[…]
Maximal wp-based performance increase was around twenty percentage points for both long-generation benchmarks and less than two percentage points for short-generation and multiple choice benchmarks. This difference in scale and the above difference in optimal PLW shows that the relationship between PLW and model performance is strongly dependent on the benchmark task.
Clear qualitative performance trends for the AlpacaData- and AlpacaDataCleaned-trained models and for any model evaluated on the six translation benchmarks could not be identified.
@@ 5.3.2 Weight Regularization
5.3.3 Data Memorization
We next explored how PLW affected training data memorization. We sampled 10,000 unique prompts from the AlpacaDataShort training set, generated completions from each prompt, and calculated corpus BLEU-4 scores. We found that for PLW from 0.0 to around 0.1, models memorized most of the training data, consistently scoring near 80 corpus BLEU. Corpus BLEU then decreased as PLW increased from 0.1. We also analyzed generation length on the AlpacaEval 1 test set, which showed a generally increasing trend with PLW.222 Note that recent work Li et al. (2023); Dubois et al. (2024) has shown that AlpacaEval 1 has a preference for long generations, and we argue that improved AlpacaEval scores are not simply due to longer generations. First, while AlpacaEval performance showed a nearly strictly-increasing relationship with PLW, generation length did not. Second, we used Mixtral as the auto-evaluator which showed a much lower length preference than the default evaluator (0.63 and 0.75, respectively). Since AlpacaDataShort is dominated by short-completion instances, we concluded that non-zero PLW decreases overfitting by allowing the model to learn generation patterns from the prompt without negatively impacting instruction-completion alignment.
5.3.4 Interpretation
Based on the above analysis, we suggest that the causal mechanisms between PLW and downstream performance of models fine-tuned on short-completion data are
  1. 1.
    preservation of PTLM weights for small PLW and
  2. 2.
    reduced overfitting (and increased generation length) for large PLW.
A tradeoff between these two mechanisms would explain the positive trend seen in the AlpacaEval and PandaLM benchmarks and the negative quadratic relationship in several of the other benchmarks.
@@ 6 Supplemental Experiments
In this section, we present supplemental experiments that suggest that PLW cannot be replaced by alternative regularization techniques for SIFT and that the effects of PLW extend to other short-completion datasets. Since the translation benchmarks showed high levels of noise and unclear correlation with PLW in the main experiment, they were not included in supplemental experiments.
@@ 6.1 Alternative Regularizers
To investigate if the effects of PLW on short-completion SIFT can be emulated with other common regularization techniques, we repeated the AlpacaDataShort training runs for PLW $=0$ and PLW $=1$ while applying various regularizers. We tested weight decay, minimizing the Minkowski distance between PTLM weights and learned weights, dropout, and label smoothing. See Appendix C.1 for visualizations.
@@ 6.2 Alternative Datasets
We repeated our main training experiment using two additional datasets, a cleaned and binarized version of UltraFeedback (Cui et al., 2023; denoted UltraFeedbackCleaned) and databricks-dolly-15k (Conover et al., 2023; denoted DatabricksDolly). We also trained on prompt-inverted versions of both datasets, denoted UltraFeedbackShort and DatabricksDollyShort, respectively, and expanded analysis to include LLaMA 3 8B AI@Meta (2024) as a PTLM. See length statistics for these four datasets in table 1 and Appendix C.2 for additional visualizations. Note that compared with the unmodified datasets used in the main experiment, both UltraFeedback and DatabricksDolly have significantly lower generation ratios of $R_{g}=1.77$ and $R_{g}=0.83$, respectively. These datasets were chosen to demonstrate the effects of PLW when fine-tuning on data with relatively balanced generation ratios.
The combined relative aggregate scores for models trained on UltraFeedbackCleaned and DatabricksDolly and for models trained on the prompt-inverted UltraFeedbackShort and DatabricksDollyShort datasets can be seen in figure 3. Visual inspection suggests that PLW affected learning for both groups of datasets. For the shortened variants, performance appears to have the same negative quadratic relationship with PLW as the AlpacaDataShort models did in the main experiment. The relationship between PLW and performance for models fine-tuned on the unmodified datasets is weaker, but appears to be generally increasing.
The clear relationship in the shortened data variants shows that the results of the main experiment extend to additional datasets. Furthermore, the weak relationship between PLW and performance of the unmodified data variants suggests that the effects of PLW are indeed dependent on the generation ratio. See Appendix D for an approach to predicting optimal PLW based on the generation ratio of the SIFT dataset.
@@ 7 Conclusion
We found that PLW had a statistically significant effect on learning for our short-completion dataset, and proper tuning of PLW allowed short-completion-trained models to outperform long-completion-trained models on short-generation benchmarks. We showed that the causal mechanism was due to a balance between two different regularizing effects and not due to increased training stability as is commonly attributed. We showed that the measured relationship extends to additional SIFT datasets and that the effects could not be sufficiently emulated with other regularizers.
@@
    Given the importance of PLW and given that many SIFT datasets and almost all natural language understanding (NLU) datasets are short-completion datasets, we warn SIFT API providers about the need for a PLW parameter to adequately cover a full range of modeling applications.
@@
    We analyzed prompt loss weighting (PLW) for instruction fine-tuning LLMs. We characterized seven fine-tuning datasets by their relative completion-prompt length ratios and reported on the effect of PLW when training on each dataset. It would be helpful to extend this research to a wider range of datasets to increase the strength of our conclusions and create more complete guidelines for prompt loss weighting.
@@
    LLM-as-evaluator approaches like PandaLM and AlpacaEval are still relatively new, and these approaches are being actively developed. We chose to use Mixtral 8x7B as an auto-evaluator for AlpacaEval 1 due to budget limitations. While we cite high human evaluation correlation with Mixtral and justify this decision in section 4.4, using the default auto-evaluator would be beneficial for better comparison with other research.
@@ Acknowledgements
This work was supported by the Technology Innovation Program funded by the Korean Ministry of Trade, Energy, and Industry (MOTIE, Korea) (No. 20014406, “Development of interactive sign language interpretation service based on artificial intelligence for the hearing impaired”) and the Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City (No. BA00000797, LLM-based sign language translation for weather forecasts).
@@ Appendix A Prompt Inversion
In order to experiment with short-completion fine-tuning datasets, we propose using prompt inversion to rotate the instruction, input, and output fields of instances. Prompt inversion modifies an instance to use the original instruction as the new completion text and the original output as the new input text. The model is then given the following instruction:
@@
Synthesis of the original input and output texts to predict the original instruction requires both language understanding and reasoning, and prompt-inverted instances should be seen as natural language understanding (NLU) tasks.
To generate short versions of the instruction datasets used in our experiments, we used prompt-inversion to modify every instance with a completion-prompt length ratio $R_{g}>1$, based on the tokenized lengths of each field (where “prompt” is the concatenation of the instruction and input fields as explained in section 2.1). Thus, given any long-completion dataset, a textually-similar short-completion dataset can be generated and used for comparison. Note that unless all instances have a generation ratio $R_{g}>1$, the resulting dataset will contain a mixture of unmodified instances and prompt-inverted instances.
@@ Appendix E Reproducibility
This section provides technical details on all experiments and benchmarks for transparency and to encourage reproduction of results. To help with reproducibility, we will also uploaded the fine-tuned models, test generation outputs, and our modified datasets to the HuggingFace Hub and can be accessed at https://huggingface.co/collections/mathewhe/plw-66fe40fac6e586d8435bd563. Note that unless specified otherwise, default parameters were used for all training and testing.
@@ E.1 Model Fine-Tuning
Model fine-tuning was performed with the Stanford Alpaca GitHub repository at https://github.com/tatsu-lab/stanford_alpaca/tree/761dc5b.
@@
AlpacaData is available from the Stanford Alpaca repository. AlpacaDataCleaned can be found at https://github.com/gururise/AlpacaDataCleaned/tree/791174f and is labeled “alpaca_data_cleaned.json”. As noted above, AlpacaDataShort can be accessed at https://huggingface.co/collections/mathewhe/plw-66fe40fac6e586d8435bd563.
@@ E.2 Model Evaluation
In an effort to match the current HuggingFace Open LLM leaderboard, we evaluated ARC Challenge, TruthfulQA-MC2, WinoGrande, and PIQA on the same EEH commit that the HuggingFace leaderboard uses: https://github.com/EleutherAI/lm-evaluation-harness/tree/b281b09 We also matched the number of shots with the number used for the HuggingFace leaderboard for ARC Challenge, TruthfulQA-MC2, and WinoGrande.
TruthfulQA-Gen and all translation tasks were evaluated using a more recent commit at https://github.com/EleutherAI/lm-evaluation-harness/tree/b93c3bc. We modified the translation tasks at this commit to include an appropriate prompt to support zero-shot translation. These changes can be seen at https://github.com/mathewhuen/plw_lm-evaluation-harness/compare/b93c3bc..1957d1a.
@@ E.5 Supplemental Experiments
The second experiment introduced two new datasets: UltraFeedbackBinarizedCleaned and DatabricksDolly. UltraFeedbackBinarizedCleaned can be found at https://huggingface.co/datasets/allenai/ultrafeedback_binarized_cleaned/tree/f304ce5. And DatabricksDolly can be found at https://huggingface.co/datasets/databricks/databricks-dolly-15k/tree/bdd27f4. The modified datasets UltraFeedbackShort and DatabricksDollyShort can be accessed at https://huggingface.co/collections/mathewhe/plw-66fe40fac6e586d8435bd563.